OpenStudy (anonymous):

For which value of x does the graph of y = 2x^2 − 7x + 6 cross the x-axis? -3/2 -2/3 2 3

OpenStudy (lgbasallote):

equate y to zero 2x^2 - 7x + 6 = 0 now factor (2x+3)(x+2) equate the factors to zero 2x + 3 = 0 AND x + 2 = 0 can you do it from there?

OpenStudy (anonymous):

Sorry, no. Can you just tell me the answer?

OpenStudy (lgbasallote):

2x + 3 = 0 isolate x 2x = -3 divide both sides by 2 to get x i am not going to give you the answer..it should be obvious from there on -_-